OpenStudy (anonymous):

THREE QUESTIONS HELP GIVING MEDALS 1.) Let u = <5, 6>, v = <-2, -6>. Find -2u + 5v. <-20, -42> <-20, 0> <0, 18> <-6, 0> 2.) Find the dot product, a · b. a = 6i + 5j, b = -5i + 4j 50 <1, 9> -10 <-30, 20> 3.) Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <7, 2>, v = <21, 6> Neither Orthogonal Parallel None of these

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

@Loser66

OpenStudy (danjs):

yeah

OpenStudy (danjs):

You can add similar dimensions for vectors, and you can multiply a vector by a constant scalar....

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (danjs):

c*<x,y> = <cx, cy>

OpenStudy (danjs):

<a,b> + <c,d> = <a+c , b+d>

OpenStudy (danjs):

those are the two rules for part 1

OpenStudy (anonymous):

i literally hv no clue how to get an answer

OpenStudy (danjs):

ok.. , here is the first step, replacing u and v with the vectors given -2u + 5v = -2*<5, 6> + 5*<-2, -6>

OpenStudy (danjs):

Now you can, by the first rule i gave, distribute the numbers into the vectors...

OpenStudy (anonymous):

is it A?

OpenStudy (danjs):

havent looked at answers , not sure

OpenStudy (danjs):

-2u + 5v = -2*<5, 6> + 5*<-2, -6> = <-10, -12> + < -10, -30>

OpenStudy (danjs):

now use rule 2 that i gave

OpenStudy (danjs):

<-10, -12> + < -10, -30> = <-10+(-10) , -12 + (-30) >

OpenStudy (loser66):

@em2000 It is not hard, please learn how to solve it rather than rely on others. Follow @DanJS and you can do whatever you like. Please, don't just guess the answer. YOu CAN do it by yourself.:)

OpenStudy (anonymous):

thank u @DanJS

OpenStudy (anonymous):

for #2 i got 50 and #3 i got neither, did u get those?

OpenStudy (danjs):

2-dimensional dot product for part 2 a = <6 , 5> b = <-5 , 4> a * b = 6*-5 + 5*4 Dot product, just add the product of each dimension

OpenStudy (anonymous):

-10?

OpenStudy (danjs):

right

OpenStudy (anonymous):

awesome, is 3 wrong?

OpenStudy (danjs):

For part 3, the notes to remember are, if for 2 vectors A and B A * B = 0 They are Orthogonal One is a scalar multiple of the other, they are parallel, B = k*A or A = k*B

OpenStudy (anonymous):

so they def not orthogonal

OpenStudy (danjs):

hint, the dot product is not zero, and try multiplying the first vector by 3

OpenStudy (anonymous):

ohhh theyre parallel!!

OpenStudy (danjs):

yep

OpenStudy (anonymous):

thank u so much you are great!!!

OpenStudy (danjs):

no prob, numbers are meaningless, just remember the 3 or 4 rules... good luck